OpenStudy (anonymous):

Find the absolute maximum and absolute minimum values of f(x)= x^3 - 6x^2 +9x -3 on interval [-1,2 ]

OpenStudy (anonymous):

f(x)= x^3 - 6x^2 +9x -3 derivate: 3x^2 - (2)x + 9 set it equal to zero 0=3x^2 - (2)x + 9 solve for x

OpenStudy (shamil98):

what he said ,lol

OpenStudy (anonymous):

i did that and got 3(x-3)(x-1) but then what

OpenStudy (anonymous):

my mistake derivative is: 3x^2 - (2)(6)x + 9 (forgot the 6)

OpenStudy (anonymous):

ya, good so what are your roots then

OpenStudy (anonymous):

well what i did was for the deriv. i got 3x^2 -12x +9. and set it equal to zero to get 3(x-3)(x-1). I believe i should set the (x-3) and (x-1) to zero but i cant remembe rwhat to do with the 3 in the front

OpenStudy (anonymous):

you don't need to worry about the 3 infront, just do, 0=x-3 0=x-1

OpenStudy (anonymous):

ohhhhh okay thank you :)

OpenStudy (anonymous):

yep ^_^